OpenStudy (anonymous):

is sinA cscA=1 true.?

OpenStudy (aravindg):

csc A=?

OpenStudy (anonymous):

1/sinu

OpenStudy (aravindg):

sin A *1/sin A =?

OpenStudy (anonymous):

1/csc

OpenStudy (aravindg):

try again?

OpenStudy (aravindg):

sin A * (1/sin A) =?

OpenStudy (anonymous):

csc

OpenStudy (aravindg):

nooo!! 1 is the answer! sin a * 1/sin A =1

OpenStudy (aravindg):

sin A cancels out.

OpenStudy (aravindg):

hence its true.

OpenStudy (anonymous):

ohh okay thanks

OpenStudy (aravindg):

yw :)